Andy Martinez/NBL: Andy will be generating the reports for the 2006 run of the RUL Old In-Process items report. The reports will be sent out towards the end of September. Andy will send out an e-mail asking everyone to check their water supplies and let him know of any that need to be ordered. A supply order proposal will be submitted in October. Systems will be withdrawing all items charged to missing user before 7/2005 in October.
Krista Dandurand/Physics: Krista is measuring the collections at Physics as part of prep work towards a shift of the collections.
L Melanie Robinson/Math: Melanie is working on scheduling and training students. She stated that it is very busy at her site.
Paul Young/Gov. Docs: At Alexander, Paul completed a shift of the DocLaw US collection. At LSM, Paul completed a shift of refcom cd's.
Roselyn Riley/Art: Roselyn reported that it has been a busy at her site.
Rob Krack/Alex: Rob reported that Alexander CM staff has taken over shelving for the East Asian collection. Rob reported on a water emergency that occurred at Alexander in August, above the third floor Stacks area. As staff were getting ready to close the library, the water leak was discovered. Facilities later reported that a drainage pipe overflowed or got clogged. Alexander staff pulled approximately 6-7 tables full of wet books to dry and where able to save them all. Plastic was laid over the shelves and is still in place. Facilities are going to have to check the drainage pipes regularly to prevent another water emergency. Andy advised staff not to leave plastic wrapped around shelving because it creates an environment for mold. It is better to drape over the top of the shelves, which allows for airflow, once the immediate water danger is passed. Alexander CM staff completed a shift of the Reference Collection. Rob reported that he has less opportunity to go into stacks because of the low number of part time employees.
Eddie Suarez/LSM: Eddie still is working on LSM weeding project, pulling items to be processed by LSM DTS. Several shelves from the Folio Collection where removed to create an oversized Folio section. Eddie reported that the mini-shift of the 3rd floor Periodical Collection would continue, now that part time employees are back to work. Eddie also reported that the A's are getting crowded. Shelving is under control and they are going to hold off on shelf reading this year. The Reference Collection needs to be put in order.
Brian Stubbs/Alex: Brian reported that there is a lot of work to be completed with taking over the CM responsibilities for the East Asian Library.
Eugene McElroy/SMLR: Eugene shifted 350 bound serials from first floor to compact shelving located in the basement, which opened up shelving space on the first floor. Patrons can page the serials that were moved to the basement.
Penny Weniger/Douglass: Penny reported on the tour that the Security & Facilities Group took of the Dana Library. As part of the tour, Dana staff reported on how they have a contract for an outside vendor to shelve and shelf read their collections.
Melody Tomaszewicz/DTS-LSM: Melody reported that LSM-DTS staff are up to date on the weeding project. They are getting close to 35, 000 items pulled from the collections, but thinks it will be closer to 40,000 once they weed items from other libraries.
Bill Lynch/Douglass & Kilmer: Brian reported completing a mini-shift in the East Room. He also reported that he is currently working a split shift between the Douglass Library and Kilmer Library.
Marc Forster of LSM asked the group what CM information they would like included in a "desktop" feature that he has developed. After discussion, the group decided that they would like the following items included on the desktop:
If possible, the group would like to see a copy of the missing book form adapted so it could be used directly from the desktop. Roselyn asked that a date line be added because the current form does not have one. Melanie volunteered to see if she could create an on-line version.
The group continued a discussion as a follow up on an agenda item from a past meeting. As a follow up, Andy asked the group for volunteers to test out whether enhanced shelf reading could work. Staff from the Art Library, Douglass Library, Math Library, and Physics Library volunteered their sites. Work will start at these sites over the next month.
The group reviewed the draft document concerning "In-Process" procedures. The document incorporated the suggestions from the previous meeting, using screen captures to detail the work. Each suggested action was discussed, with changes and corrections noted. The document will be revised again and forwarded to the group for review.
Date: Wednesday, October 11, 2006
Time: 2:30 pm-4: 30 pm
Primary Location: Media Center Classroom, Kilmer Library, and Livingston Campus
Satellite Location(s): Dana Library - Electronic Classroom II
Camden Library - Conference Room 290